solve f(x)=-1/2x²+3x-5/4
Answer by stanbon(75887)

f(x) = -1/2x²+3x-5/4
----
a = -1/2 ; b = 3 ; c = -5/4
----
Quadratic Formula:
x = [-3 +- sqrt(9 - 4*(-1/2)(-5/4))]/-1
------
x = [-3 +- sqrt(9-2.5)]/-1
---
x = [-3 +- sqrt(6.5)]/-1
----
x = (-3 +2.55]/-1 = +0.45
OR
x = (-3 - 2.55)/-1 = +5.55
===================
Cheers,
Stan H.
===================